OpenStudy (anonymous):

Given that the curve y=ax^2 + 3ax + a + b, where a≠0, intersects the x-axis at exactly one point, find b in terms of a.

OpenStudy (anonymous):

The answer given is 5/4 a

OpenStudy (anonymous):

But when I tried working it out I got 1/4 a

OpenStudy (anonymous):

\[a= a, b=3a,c=a+b\]\[3a ^{2}-4(a)(a+b)=0\]\[3a ^{2}-4a ^{2}+4ab=0\]\[-a ^{2}+4ab=0\]\[4ab=a ^{2}\]\[4b=a\]\[b=\frac{ 1 }{ 4 }a\]

OpenStudy (anonymous):

This is what I did

OpenStudy (anonymous):

erm ill try and help let me have a look.

OpenStudy (anonymous):

alright thanks!

OpenStudy (anonymous):

@JamesWolf

OpenStudy (anonymous):

trying to solve it with the quadratic equation, which is not your method gives me -3/2, so I'm abit stuck really, sorry

OpenStudy (anonymous):

oh alright then.... Thanks anyways.

OpenStudy (anonymous):

Im annoyed. I dont understand why it didnt work

OpenStudy (anonymous):

if it only crosses the x axis once, then that means it looks like this 

OpenStudy (dumbcow):

i got it (3a)^2 = 9a^2 not 3a^2

OpenStudy (dumbcow):

other than that @BelleFlower your steps are correct

OpenStudy (anonymous):

or some flip of that. so if you use the quadratic equation (-b +- sqrt(b^2 -4ac))/2a the square root, sqrt(), should equal 0. since there is only one x intercept. So were left with -3a/2a = -3/2 @dumbcow, where have I gone wrong here?